OpenStudy (anonymous):

Another medal if answered A right circular cone has a height of 3.75 in. The diameter of the base is 1.2 in. Find the volume of the cone.

OpenStudy (anonymous):

use the cylender volume formula!

OpenStudy (anonymous):

1/3 H \[Pi] r^2

OpenStudy (anonymous):

Integrate[\[Pi] (r*h/H)^2, {h, 0, H}] --> 1/3 H \[Pi] r^2

OpenStudy (farmdawgnation):

ktklown - you can use the equation editor to make that more readable. :)

OpenStudy (anonymous):

yeah, that takes more effort :P

OpenStudy (farmdawgnation):

LOL. Nice. \[\frac{1}{3}\pi r^2\] Is that what you were going for? :P

OpenStudy (anonymous):

Yeah.... if they would render mathematica expressions i'd use it ;-)

OpenStudy (farmdawgnation):

Ah forgot the height \[\frac{1}{3} H \pi r^2\]

OpenStudy (anonymous):

I need the full answer in inches

OpenStudy (farmdawgnation):

Well, it's a really simple equation to solve. Just plug these values into that formula above: \[H = 3.75\] \[r = 0.6\] Then just run the equation through a calculator and you're done. :) Also worth noting the 0.6 for r is because the equation asks for the radius. Which is half of the diameter, which the problem tells you is 1.2.